OpenStudy (anonymous):

Gethan put $4000 in a 2-year CD paying 5% interest, compounded monthly. After 2 years, he withdrew all his money. What was the amount of the withdraw?

OpenStudy (kropot72):

\[A _{t}=P(1+\frac{r}{12})^{2\times 12}\] where At is the amount after 2 years, P is the principal and r is the annual interest as a decimal. \[A _{t}=4000(1+\frac{0.05}{12})^{24}=you\ can\ calculate?\]